a tower subtends angle z, 2z ,3z respectively at points A,Band C all lying on a horizontal line through the foot of the tower , then AB/BC is ?

Grade:12th pass

1 Answers

ng29
8 years ago
its just a labour questions we had to solve three equations to get the ratio.
let the ht of tower be h  and  O be bottom most point of tower
tan(3z) = h/OC
tan(2z) = h/(OC+BC)
tan(3z) = h/(OC+BC+AB)
SOLVE THESE EQ TO GET THE ANSWER (trigonometry can be used)
approve if useful